Graph the linear system and estimate the solution:
3x+y=9
x-2y=10

Answer by ewatrrr(24785)


Hi
Graphing the linear system and estimate the solution
3x+y=9 OR y = -3x + 9 Pt(3,0), Pt(1,6), pt(2,3)
x-2y=10 OR y = x/2 - 5 Pt(0,-5), Pt(2, -4), Pt(4,-3)
solution for the system is the point of intersection of the lines